A chemist reacted 17.25 grams of sodium metal with an excess amount of chlorine gas. The chemical reaction that occurred is shown.


__Na + __ Cl2 → __ NaCl


If the percentage yield of the reaction is 88%, what is the actual yield? Show your work, including the use of stoichiometric calculations and conversion factors. 


1
Expert's answer
2021-04-07T11:29:00-0400

Given,

% yield ="\dfrac{\text{ Reacted Na }}{\text{ Required Na}}\times 10"


"88=\dfrac{17.25}{\text{ Required (Na)}}\times 100"


"\text{ Required Na}=\dfrac{1725}{88}=19.60g"


Actual yield = "\dfrac{\text{ Reacted Na }}{\text{ Required Na}}\times 100"


"=\dfrac{17.25}{19.60}\times 100=88%" %